# Control Barrier Functions via Minkowski Operations for Safe Navigation among Polytopic Sets

### v2

@@ IV-B CBF Construction Using Signed Distance
The point $\bm{z}^{*}$, defined in $\mathcal{M}_{\mathcal{D}}$-space and obtained from our optimization-based signed distance formulation, is referred to as the critical point, as its derivative is used in computing the derivative of CBF. Geometrically, $\bm{z}^{*}$ lies on the boundary of the CO, $\partial\mathcal{O}^{\mathcal{C}}$, and is the closest point to the origin $\bm{0}_{c}$ in $\mathcal{M}_{\mathcal{D}}$-space. Thus, the signed distance is equal to $\|\bm{z}^{*}\|_{2}$.
@@
$\begin{split}&h(\bm{x})=\text{sd}(\mathcal{R}(\bm{x}),\mathcal{O})-d_{\text{safe}}=\text{sd}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}}(\bm{x}))-d_{\text{safe}},\\[3.00003pt] &\,\,\text{where}\,\,\text{sd}(\mathcal{R}(\bm{x}),\mathcal{O})=\left\{\begin{matrix}\|\bm{z}^{*}(\bm{x})\|_{2}\,\,\,\,\,\quad\text{if}\,\,\,\bm{0}_{c}\notin\mathcal{O}^{\mathcal{C}}\\ -\|\bm{z}^{*}(\bm{x})\|_{2}\quad\text{if}\,\,\,\bm{0}_{c}\in\mathcal{O}^{\mathcal{C}}\end{matrix}\right.,\end{split}$ (14)
where $\bm{z}^{*}(\bm{x})\in\mathbb{R}^{d}$ is the optimal solution to (10) if $\mathcal{R}\cap\mathcal{O}=\varnothing$, and can be indirectly found by solving (12) with its dual variable analysis in the case of $\mathcal{R}\cap\mathcal{O}\neq\varnothing$. Recall that both the critical point $\bm{z}^{*}(\bm{x})$ and the CO, $\mathcal{O}^{\mathcal{C}}(\bm{x})$, depend on the robot’s state $\bm{x}$. The associated safe set of the system is defined as $C=\{\bm{x}\in\mathbb{R}^{n}:\text{sd}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}}(\bm{x}))\geq d_{\text{safe}}\}.$ Note that the SDF-based CBF (14) is implicit and is obtained from optimization that depends on system state $\bm{x}$. It is important to note the SDF in Euclidean space is differentiable almost everywhere (a.e.) [20], which means it is differentiable at all points except for a set of (Lebesgue) measure zero.
@@
The optimal solution $\bm{z}^{*}(\bm{x}(t))$ of (10) is unique, continuous, and differentiable a.e. for all $[t_{0},T]$.
@@
    By Lem. 2, the CO set $\mathcal{O}^{\mathcal{C}}(\bm{x})$ is always closed and convex, and therefore, the Euclidean projection of the origin $\bm{0}_{c}$ on the set $\mathcal{O}^{\mathcal{C}}(\bm{x})$ is unique [39, Chap. 8]. As (10) is a strictly convex QP, its solution $\bm{z}^{*}(\bm{x}(t))$ is continuous and differentiable a.e based on [41, Thm. 1]. ∎
@@
In the penetration depth case, our optimization formulation empirically provides accurate derivative estimates. However, a formal nonsmooth analysis, similar to that in [18, 42], along with a derivation of the corresponding derivatives, is left for future work. Importantly, this does not affect our derivation for the minimum distance case, as these two events are complementary. In the next subsection, we focus specifically on the minimum distance case.
@@ IV-C Derivative Calculation Using Differentiable Optimization
To enforce CBF constraint to system dynamics, we need to find the derivative of $h(\bm{x})$. However, the CBF $h(\bm{x})$ (14) depends on the solution to an optimization problem $\bm{z}(\bm{x})$, defined implicitly to state $\bm{x}$. Therefore, we cannot differentiate it w.r.t. $\bm{x}$ directly. Given this implicit dependence, we first apply the chain rule to express $\partial h/\partial\bm{x}$ as:
@@
where $\frac{\partial h}{\partial\bm{z}^{*}}=\frac{\bm{z}{{}^{*}}^{\top}}{|h|}\in\mathbb{R}^{1\times d}$. The second term $\frac{\partial\bm{z}^{*}}{\partial\bm{x}}\in\mathbb{R}^{d\times n}$ is the Jacobian of the optimal solution $\bm{z}^{*}$ to (10), which describes how it changes as inequality constraints of CO vary with the state $\bm{x}$.
Proposition 1.
Let $\bm{z}^{*}(\bm{x}(t))$ denote the optimal solution to the QP (10). Then its Clarke generalized subdifferential [43] $\partial\bm{z}^{*}(\bm{x}(t))$ is nonempty for all $[t_{0},T]$, and has a single unique Jacobian for all but a measure zero set of points $\bm{x}(t)$.
@@
    This follows directly from [41, Thm. 1], as the objective function of QP (10) is strictly convex and the inequality constraints depend smoothly on $\bm{x}(t)$ by Assumption 2). ∎
@@
$\begin{split}G(\bm{z},\bm{\lambda})=\begin{bmatrix}2I_{d}\bm{z}+({A^{\mathcal{C}}})^{\top}\bm{\lambda}\\ D(\bm{\lambda})(A^{\mathcal{C}}\bm{z}-b^{\mathcal{C}})\end{bmatrix},\end{split}$
@@
$\begin{split}\partial_{\bm{\xi}}G(\bm{\xi}^{*})&=\begin{bmatrix}2I_{d}&({A^{\mathcal{C}}})^{\top}\\ D(\bm{\lambda}^{*})A^{\mathcal{C}}&D(A^{\mathcal{C}}\bm{z}^{*}-b^{\mathcal{C}})\end{bmatrix},\\[4.30554pt] \partial_{\bm{x}}G(\bm{\xi}^{*})&=\begin{bmatrix}d_{\bm{x}}({A^{\mathcal{C}}})^{\top}\bm{\lambda}^{*}\\ D(\bm{\lambda}^{*})\left(d_{\bm{x}}A^{\mathcal{C}}\bm{z}^{*}-d_{\bm{x}}b^{\mathcal{C}}\right)\end{bmatrix},\end{split}$ (17)
@@
Since (17) is derived by differentiating the KKT conditions, complementary slackness always holds. Thus, it is sufficient to consider only the active constraints when computing these derivatives. Specifically, we can replace $A^{\mathcal{C}},b^{\mathcal{C}}$ with $A^{\mathcal{C}}_{k_{\text{act}}},b^{\mathcal{C}}_{k_{\text{act}}}$.
@@ IV-D Configuration Obstacle Computation
Consider a polygonal robot modeled as a rigid body in SE(2), with configuration (position and orientation) is $\bm{q}=[x,y,\theta]^{\top}$, and a nonzero initial configuration $\bm{q}=[x_{0},y_{0},\theta_{0}]^{\top}$. Its shape in H-rep can be parameterized by its configuration in $\mathcal{W}$-space as $\mathcal{R}(\bm{q})=\{y\mid A(\bm{q})y\leq b(\bm{q})\}$, where $A(\bm{q})=AR(\theta-\theta_{0})^{\top},\leavevmode\nobreak\ b(\bm{q})=b+A(\bm{q})\bm{p}-A\bm{p}_{0}$ with $\bm{p}=[x,y]^{\top}$, $\bm{p}_{0}=[x_{0},y_{0}]^{\top}$, and $A,b$ defining the robot’s shape at the initial configuration. $R(\theta)=[\cos\theta,-\sin\theta;\sin\theta,\cos\theta]$ is a rotation matrix.
@@
In contrast, when rotation is involved, both the slope and order of the bounding hyperplanes change, leading to nontrivial deformations of the CO. While Minkowski sum can be computed via linear projection [45], this method also does not provide an analytical solution for the CO. Deriving an explicit relationship between the robot and obstacle representations in $\mathcal{W}$-space and the CO in $\mathcal{M}_{\mathcal{D}}$-space remains an open problem for future research.
@@ IV-E Controller Synthesis Using CLF-CBF-QP
We discretize time and hold the state fixed at the beginning of each interval, rendering the constraints affine in the control input and casting the problem as a QP. We use the standard CLF-CBF-QP (18) introduced in [2] to compute the optimal control $\bm{u}(t)$, which is applied to update the dynamics (1).
[…]
@@ V Case Study and Implementation
Figure 2: Top-Left: The robot’s trajectory (green) moving left-to-right from the start (${t=0}$) to the goal (${t=8}$). Snapshots show the robot (blue) avoiding the obstacle (red). Top-Right: The corresponding motion of the CO (green) moving right-to-left from ${t=0}$ to ${t=8}$ in $\mathcal{M}_{\mathcal{D}}$-space  (see video https://youtu.be/3Dh0gtDW8bE). Bottom: Control input $\bm{u}(t)$, and CBF $h(\bm{x})$ profiles. Both control inputs satisfy actuator limits and safety constraints
@@ V-A Case I: Single-Integrator Model (Pure Translation)
We first consider the scenario of a triangular robot $\mathcal{R}$ with single-integral dynamics translating to the desired position $\bm{x}_{d}$ while avoiding a polygonal obstacle $\mathcal{O}$: $\bm{x}=[x,y]^{\top},\,\,\dot{\bm{x}}=\bm{u}=[u_{1},u_{2}]^{\top}$, where $x,y$ denote the position, and $u_{1},u_{2}$ are the control inputs (linear velocites). The dynamics can be rewritten in form of (1) with $f(\bm{x})=\bm{0}_{2\times 1}$ and $g(\bm{x})=\bm{I}_{2\times 2}$. We use the proposed signed distance CBF (14) to impose safety and a CLF $V(\bm{x})=(\bm{x}-\bm{x}_{d})^{2}$ to reach $\bm{x}_{d}=(5,8)$. The other parameters are $\bm{x}(0)=(-4.83,0.77),\bm{u}_{\max}=-\bm{u}_{\min}=5$m/s, $d_{\text{safe}}=0,p=10,c=1,\gamma=3$, $\epsilon=0$.
Figure 3: Left: Four columns depict movement over time. At $t=0$ The robot starts in collision with the obstacle in $\mathcal{W}$-space (top), and the corresponding CO contains the origin in $\mathcal{M}_{\mathcal{D}}$-space  (bottom). As the robot moves out of the obstacle, the CO moves away from the origin ($t=0$ to $t=2.15$ top and bottom). The collision is resolved by time $t=2.15$ seconds). Right-Top: Snapshots of robot motion in $\mathcal{W}$-space demonstrate the non-conversativeness of the proposed framework. Right-Bottom: The CBFs $h_{i}(\bm{x})$ for each obstacle stay above the safety margin over time, thereby guaranteeing safety.
The derivative of the CBF is computed using (16) and (17). Each control loop involves computing the CO (6) and solving a QP for minimum distance (10) and optimal control (18). Loops run at 100 Hz and take 7.38 ms on average. The simulation results (Fig. 2) demonstrate the non-conservativeness of the proposed CBF, i.e., the robot safely moves in close proximity to the obstacle boundary.
@@ V-B Case II: Unicycle Model
We highlight that the CO depends on both position and orientation, unlike $\mathcal{W}$-space obstacles, which depend only on position. Thus, high-order CBF [4] or auxiliary-variable CBF [5] is unnecessary to handle the high relative degree constraint. We use two CLFs $V_{1}(\bm{x})=(\theta-\tan^{-1}(\frac{y_{d}-y}{x_{d}-x}))^{2}$, $V_{2}(\bm{x})=(v-v_{d})^{2}$ [4] to reach the goal.
@@ V-B2 Unicycle Collision Avoidance with Multiple Obstacles
For the multiple obstacles scenario, the safety is guaranteed by ensuring that safety constraint is satisfied w.r.t. each Obstacle $\mathcal{O}_{i},i\in\{1,\ldots,4\}$, as mentioned in Rem. 2, i.e., $h_{i}(\bm{x})=\text{sd}(\mathcal{R}(\bm{x}),\mathcal{O}_{i})-d_{\text{safe}}=\text{sd}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}}_{i}(\bm{x}))-d_{\text{safe}}\geq 0.$ The associated CBF constraints are enforced in the CLF-CBF-QP as $L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\bm{u}+\gamma h_{i}(\bm{x})\geq\epsilon$ The parameters used are the same as in the previous setting, except for $\bm{x}_{d}=(3,13.5),v_{d}=1.8$m/s, $p=10,c=6,\gamma=0.3,\epsilon=0$.

### v3

@@ IV-B CBF Construction Using Signed Distance
The point $\bm{z}^{*}$, defined in $\mathcal{M}_{\mathcal{D}}$-space and obtained from our optimization-based signed distance formulation, is referred to as the critical point, as its derivative is used in computing the derivative of CBF. Geometrically, $\bm{z}^{*}$ lies on the boundary of the CO, $\partial\mathcal{O}^{\mathcal{C}}$, and is the closest point to the origin $\bm{0}_{c}$ in $\mathcal{M}_{\mathcal{D}}$-space. Thus, the signed distance is equal to $\pm\|\bm{z}^{*}\|_{2}$.
@@
[…]
where $\bm{z}^{*}(\bm{x})\in\mathbb{R}^{d}$ is the optimal solution to (10) if $\mathcal{R}\cap\mathcal{O}=\varnothing$, and can be indirectly found by solving (12) with its dual variable analysis in the case of $\mathcal{R}\cap\mathcal{O}\neq\varnothing$. Recall that both the critical point $\bm{z}^{*}(\bm{x})$ and the CO, $\mathcal{O}^{\mathcal{C}}(\bm{x})$, depend on the robot’s state $\bm{x}$. The associated safe set of the system is defined as $C=\{\bm{x}\in\mathbb{R}^{n}:\text{sd}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}}(\bm{x}))\geq d_{\text{safe}}\}.$ Note that the SDF-based CBF (14) is implicit and is obtained from an optimization that depends on system state $\bm{x}$. It is important to note that the SDF in Euclidean space is differentiable a.e. [20], which means it is differentiable at all points except for a set of (Lebesgue) measure zero.
@@
Let $\mathcal{X}_{f}=\{x\in\mathcal{X}\mid\mathcal{R}(\bm{x})\cap\mathcal{O}=\varnothing\}$. The optimal solution $\bm{z}^{*}(\bm{x})$ of (10) is unique for every $\bm{x}\in\mathcal{X}_{f}$, and $\bm{z}^{*}(\cdot)$ is continuous and differentiable a.e. on $\mathcal{X}_{f}$.
@@
    By Lem. 2, the CO set $\mathcal{O}^{\mathcal{C}}(\bm{x})$ is nonempty, closed, and convex, and therefore, the Euclidean projection of the origin $\bm{0}_{c}$ on the set $\mathcal{O}^{\mathcal{C}}(\bm{x})$ is unique [40, Chap. 8]. As (10) is a strictly convex QP, its solution $\bm{z}^{*}(\bm{x})$ is continuous and differentiable a.e. based on [42, Thm. 1]. ∎
@@
Extending Lemma 3 to the penetration depth case is deferred to future work. The min-distance case is a strictly convex QP, while penetration depth is an LP, so the differentiability results of [42] do not directly apply. The min-distance and penetration depth cases are complementary, and we focus on the min-distance case in the next subsection. While our LP formulation performs well empirically and yields stable derivative estimates under the same $\mathcal{M}_{\mathcal{D}}$-space framework, a rigorous differentiability analysis of the penetration depth case is planned for future work.
@@ IV-C Derivative Calculation Using Differentiable Optimization
To enforce the CBF constraint on system dynamics, we need to find the derivative of $h(\bm{x})$. However, the CBF $h(\bm{x})$ (14) depends on the solution to an optimization problem $\bm{z}(\bm{x})$, defined implicitly to state $\bm{x}$. Therefore, we cannot differentiate it w.r.t. $\bm{x}$ directly. Given this implicit dependence, we first apply the chain rule to express $\partial h/\partial\bm{x}$ as:
@@
where $\partial h/\partial\bm{z}^{*}=\bm{z}{{}^{*}}^{\top}/(h+d_{\text{safe}})\in\mathbb{R}^{1\times d}$. The second term $\partial\bm{z}^{*}/\partial\bm{x}\in\mathbb{R}^{d\times n}$ is the Jacobian of the optimal solution $\bm{z}^{*}$ to (10), which describes how it changes as inequality constraints of CO vary with the state $\bm{x}$.
Corollary 1.
Let $\bm{z}^{*}(\bm{x})$ be the optimal solution to the QP (10) for $\bm{x}\in\mathcal{X}_{f}$. Its Clarke generalized subdifferential [43] $\partial\bm{z}^{*}(\bm{x})$ is nonempty for all $\bm{x}\in\mathcal{X}_{f}$, and has a single unique Jacobian for all but a measure zero set of points.
@@
    This follows directly from [42, Thm. 1], as the objective function of QP (10) is strictly convex. ∎
@@
[…]
@@
[…]
@@
As (17) is derived by differentiating the KKT conditions; complementary slackness holds. Thus, it is sufficient to consider only the active constraints when computing the derivatives. That is, we can replace $A^{\mathcal{C}},b^{\mathcal{C}}$ with $A^{\mathcal{C}}_{k_{\text{act}}},b^{\mathcal{C}}_{k_{\text{act}}}$.
@@ IV-D Configuration Obstacle Computation
Consider a polygonal robot modeled as a rigid body in SE(2), with states (position and orientation) is $\bm{x}=[x,y,\theta]^{\top}$, and a nonzero initial configuration $\bm{x}_{0}=[x_{0},y_{0},\theta_{0}]^{\top}$. Its shape in H-rep can be parameterized by its configuration in $\mathcal{W}$-space as $\mathcal{R}(\bm{x})=\{y\mid A(\bm{x})y\leq b(\bm{x})\}$, where $A(\bm{x})=AR(\theta-\theta_{0})^{\top},\leavevmode\nobreak\ b(\bm{x})=b+A(\bm{x})\bm{p}-A\bm{p}_{0}$ with $\bm{p}=[x,y]^{\top}$, $\bm{p}_{0}=[x_{0},y_{0}]^{\top}$, and $A(\bm{x}_{0}),b(\bm{x}_{0})$ defining the robot’s shape at the initial configuration. $R(\theta)=[\cos\theta,-\sin\theta;\sin\theta,\cos\theta]$ is a rotation matrix.
@@
In contrast, with rotation the bounding hyperplanes change both slope and order, causing nontrivial CO deformations. While the Minkowski sum can be computed via linear projection [45], this method still does not provide an analytical expression for CO. Deriving an explicit relationship between the robot-obstacle representations in $\mathcal{W}$-space and the CO in $\mathcal{M}_{\mathcal{D}}$-space remains an open problem for future research.
@@ IV-E Controller Synthesis Using CLF-CBF-QP
We implement (18) by discretizing the horizon $[0,T]$ into equal intervals of $\Delta t$, holding the state fixed at the start of each step. At time step $t=k\Delta t$, the control is held constant over $t_{k},t_{k+1})$, rendering the constraints affine in $\bm{u}$. We then solve the standard CLF-CBF-QP [2] with $\bm{x}(t_{k})$ to compute the optimal control $\bm{u}(t_{k})=\bm{u}_{k}$, and apply $\bm{u}_{k}$ on $t_{k},t_{k+1})$ to propagate the dynamics (1).
$\begin{split}\underset{\bm{u}(t),\delta(t)}{\min}\quad&\int_{0}^{T}\bm{u}(t)^{\top}\bm{u}(t)+p\delta(t)^{2}\\[1.99997pt] \text{subject to}\quad&L_{f}h(\bm{x})+L_{g}h(\bm{x})\bm{u}+\gamma h(\bm{x})\geq\epsilon,\\ &L_{f}V(\bm{x})+L_{g}V(\bm{x})\bm{u}+cV(\bm{x})\leq\delta,\\ &\bm{u}_{\min}\leq\bm{u}\leq\bm{u}_{\max},\end{split}$ (18)
@@ V Case Study and Implementation
Figure 2: Top-Left: The robot’s trajectory (green) moving left-to-right from the start (t=0{t=0}) to the goal (t=8{t=8}). Snapshots show the robot (blue) avoiding the obstacle (red). Top-Right: The corresponding motion of the CO (green) moving right-to-left from t=0{t=0} to t=8{t=8} in $\mathcal{M}_{\mathcal{D}}$-space  (see video https://youtu.be/3Dh0gtDW8bE). Bottom: Control input $\bm{u}(t)$, and CBF $h(\bm{x})$ profiles. Both control inputs satisfy actuator limits and safety constraints.
@@ V-A Case I: Single-Integrator Model (Pure Translation)
We first consider the scenario of a triangular robot $\mathcal{R}$ with single-integrator dynamics translating to the desired position $\bm{x}_{d}$ while avoiding a polygonal obstacle $\mathcal{O}$: $\bm{x}=[x,y]^{\top},\,\,\dot{\bm{x}}=\bm{u}=[u_{1},u_{2}]^{\top}$, where $x,y$ denote the position, and $u_{1},u_{2}$ are the control inputs (linear velocities). The dynamics can be rewritten in form of (1) with $f(\bm{x})=\bm{0}_{2\times 1}$ and $g(\bm{x})=\bm{I}_{2}$. We use the proposed SDF-CBF (14) to impose safety and a CLF $V(\bm{x})=(\bm{x}-\bm{x}_{d})^{2}$ to reach $\bm{x}_{d}=(5,8)$. The other parameters are $\bm{x}(0)=(-4.83,0.77),\bm{u}_{\max}=-\bm{u}_{\min}=5$m/s, $d_{\text{safe}}=0,p=10,c=1,\gamma=3$, $\epsilon=0$. The derivative of the CBF is computed using (16) and (17). Each control loop involves computing the CO (6) and solving a QP for minimum distance (10) and CLF-CBF-QP (18). Loops run at 100 Hz and take 7.38 ms on average. The simulation results (Fig. 2) demonstrate the non-conservativeness of the proposed CBF, i.e., the robot safely moves in close proximity to the obstacle boundary.
Figure 3: Left: Four columns depict movement over time. At $t=0$ The robot starts in collision with the obstacle in $\mathcal{W}$-space (top), and the corresponding CO contains the origin in $\mathcal{M}_{\mathcal{D}}$-space  (bottom). As the robot moves out of the obstacle, the CO moves away from the origin ($t=0$ to $t=2.15$ top and bottom). The collision is resolved by time $t=2.15$ seconds). Right-Top: Snapshots of robot motion in $\mathcal{W}$-space demonstrate the non-conservativeness of the proposed framework. Right-Bottom: The CBFs $h_{i}(\bm{x})$ for each obstacle stay above the safety margin over time, thereby guaranteeing safety.
@@ V-B Case II: Unicycle Model
We highlight that the CO depends on both position and orientation, unlike $\mathcal{W}$-space obstacles, which depend only on position. Thus, high-order CBF [4] or auxiliary-variable CBF [5] is unnecessary to handle the high relative degree constraint. We use two CLFs $V_{1}(\bm{x})=(\theta-\tan^{-1}(\frac{y_{d}-y}{x_{d}-x}))^{2}$, $V_{2}(\bm{x})=(v-v_{d})^{2}$ (similar to [4]) to reach the goal.
@@ V-B2 Unicycle Collision Avoidance with Multiple Obstacles
For the multiple obstacles scenario, the safety is guaranteed by ensuring that safety constraint is satisfied w.r.t. each Obstacle $\mathcal{O}_{i},i\in\{1,\ldots,4\}$, as mentioned in Rem. 3, i.e., $h_{i}(\bm{x})=\text{sd}(\mathcal{R}(\bm{x}),\mathcal{O}_{i})-d_{\text{safe}}=\text{sd}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}}_{i}(\bm{x}))-d_{\text{safe}}\geq 0.$ The associated CBF constraints are enforced in the CLF-CBF-QP as $L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\bm{u}+\gamma h_{i}(\bm{x})\geq\epsilon$ The parameters used are the same as in the previous setting, except for $\bm{x}_{d}=(3,13.5),v_{d}=1.8$m/s, $p=10,c=6,\gamma=0.3,\epsilon=0$.